# UniVoice: Unifying Autoregressive ASR and Flow-Matching based TTS with Large Language Models

### v1

@@ Conclusion
  • VALL-E (Wang et al. 2023). It use an autoregressive and an additional non-autoregressive model for discrete speech codec token generation.
  • NaturalSpeech2 (Shen et al. 2023). It use a non autoregressive model for continuous vectors generation.
  • F5-TTS [^2] https://huggingface.co/SWivid/F5-TTS (Chen et al. 2024c). It is a fully non-autoregressive text-to-speech system based on flow matching with Diffusion Transformer, which is trained on Emilia (He et al. 2024) dataset with around 100K hours Chinese and English speech dataset. We use the official checkpoint of F5-TTS for evaluation.
  • CosyVoice [^3] https://github.com/FunAudioLLM/CosyVoice (Du et al. 2024). It is a two-stage large-scale TTS system, first for autoregressive text-to-token generation, then a flow matching diffusion model for Mel-spectrogram generation. The model is trained on 170K hours of multilingual speech data. We use the official checkpoint of CosyVoice for evaluation.
  • FireRedTTS [^4] https://github.com/FireRedTeam/FireRedTTS (Guo et al. 2024a). It is a foundation TTS framework for industry-level generative speech applications. It includes an autoregressive text-to-semantic token model and a token-to waveform generation model. The system is trained with 248K hours of labeled speech data. We use the official pre-trained checkpoint to evaluate.
  • MaskGCT [^5] https://huggingface.co/amphion/MaskGCT (Wang et al. 2024b). It is large-scale non autoregressive TTS model without precise alignment information between text and speech following the mask-and-predict learning paradigm. It is trained on Emilia (He et al. 2024) dataset with around 100K hours Chinese and English speech dataset.
@@
  • Whisper (Radford et al. 2023). It is an advanced ASR model developed by OpenAI, designed to transcribe speech with high accuracy. It is trained on a massive dataset of 680,000 hours of supervised audio data, covering a wide range of languages and acoustic conditions. We use Whisper-small [^6] https://huggingface.co/openai/whisper-small, Whisper-large-v2777https://huggingface.co/openai/whisper-large-v2 and Whisper-large-v3888https://huggingface.co/openai/whisper-large-v3 for evaluation.
  • Paraformer (Gao et al. 2022). Paraformer is a non-autoregressive end-to-end ASR model that achieves fast parallel decoding via CIF-based acoustic modeling. It is trained on 20K hours of English data.
  • Zipformer (Yao et al. 2023). It is a faster, more memory efficient, and better-performing transformer model for speech recognition. It is trained on 960 hours LibriSpeech dataset.
@@
  • SpeechT5 (Ao et al. 2021). It is a universal pre-trained encoder-decoder model for various speech processing tasks. After pretraining, it should be fine-tuned via the loss of the downstream tasks.
  • LauraGPT (Du et al. 2023). It is a novel unified GPT-based audio LLM for audio recognition, understanding, and generation by using continuous representations for audio input and generating output audio from audio codec tokens.
  • OpusLM (Tian et al. 2025). It is a family of scalable decoder-only transformers (135M–7B parameters) that unify speech-text processing through multistream discrete token generation for both speech and text modalities. We compare the 360M and 7B models with our UniVoice.

### v2

@@ Conclusion
  • VALL-E (wang2023neural). It use an autoregressive and an additional non-autoregressive model for discrete speech codec token generation.
  • NaturalSpeech2 (shen2023naturalspeech). It use a non autoregressive model for continuous vectors generation.
  • F5-TTS [^2] https://huggingface.co/SWivid/F5-TTS (chen2024f5). It is a fully non-autoregressive text-to-speech system based on flow matching with Diffusion Transformer, which is trained on Emilia (he2024emilia) dataset with around 100K hours Chinese and English speech dataset. We use the official checkpoint of F5-TTS for evaluation.
  • CosyVoice [^3] https://github.com/FunAudioLLM/CosyVoice (du2024cosyvoice). It is a two-stage large-scale TTS system, first for autoregressive text-to-token generation, then a flow matching diffusion model for Mel-spectrogram generation. The model is trained on 170K hours of multilingual speech data. We use the official checkpoint of CosyVoice for evaluation.
  • FireRedTTS [^4] https://github.com/FireRedTeam/FireRedTTS (guo2024fireredtts). It is a foundation TTS framework for industry-level generative speech applications. It includes an autoregressive text-to-semantic token model and a token-to waveform generation model. The system is trained with 248K hours of labeled speech data. We use the official pre-trained checkpoint to evaluate.
  • MaskGCT [^5] https://huggingface.co/amphion/MaskGCT (wang2024maskgct). It is large-scale non autoregressive TTS model without precise alignment information between text and speech following the mask-and-predict learning paradigm. It is trained on Emilia (he2024emilia) dataset with around 100K hours Chinese and English speech dataset.
@@
  • Whisper (radford2023robust). It is an advanced ASR model developed by OpenAI, designed to transcribe speech with high accuracy. It is trained on a massive dataset of 680,000 hours of supervised audio data, covering a wide range of languages and acoustic conditions. We use Whisper-small [^6] https://huggingface.co/openai/whisper-small, Whisper-large-v2777https://huggingface.co/openai/whisper-large-v2 and Whisper-large-v3888https://huggingface.co/openai/whisper-large-v3 for evaluation.
  • Paraformer (gao2022paraformer). Paraformer is a non-autoregressive end-to-end ASR model that achieves fast parallel decoding via CIF-based acoustic modeling. It is trained on 20K hours of English data.
  • Zipformer (yao2023zipformer). It is a faster, more memory efficient, and better-performing transformer model for speech recognition. It is trained on 960 hours LibriSpeech dataset.
@@
  • SpeechT5 (ao2021speecht5). It is a universal pre-trained encoder-decoder model for various speech processing tasks. After pretraining, it should be fine-tuned via the loss of the downstream tasks.
  • LauraGPT (du2023lauragpt). It is a novel unified GPT-based audio LLM for audio recognition, understanding, and generation by using continuous representations for audio input and generating output audio from audio codec tokens.
  • OpusLM (tian2025opuslm). It is a family of scalable decoder-only transformers (135M–7B parameters) that unify speech-text processing through multistream discrete token generation for both speech and text modalities. We compare the 360M and 7B models with our UniVoice.